x^2-6x + ____ = 3 + ____
(x-3)^2 = ____

Answer by stanbon(75887)

x^2-6x + ____ = 3 + ____
---
x^2 - 6x + 9 = (3+(-x))^2
-------------------------------
(x-3)^2 = x^2 - 6x + 9
____
Cheers,
Stan H.
----------------